OpenStudy (anonymous):

Which statement best describes how to determine whether f(x) = x4 – x3 is an even function?

OpenStudy (perl):

Definition: (yay, i love definitions) if f(-x) = f(x), then f is even. if f (-x) = -1*f(x) , then f is odd

OpenStudy (perl):

for all x in the domain

OpenStudy (anonymous):

thanks

OpenStudy (perl):

so plug in f(-x) = (-x)^4 - (-x)^3

OpenStudy (perl):

f(-x) = (-x)^4 - (-x)^3 = x^4 - (-x)^3 = x^4 + x^3 this is neither f(x) , nor is it -f(x)